OpenStudy (anonymous):

Evaluate the limit and justify each step by indicating the appropriate limit laws lim (1 + cube root x) (2 - 6x^2 + x^3) as x approaches 8

OpenStudy (campbell_st):

substitute x = 8 and evaluate

OpenStudy (anonymous):

What about to justify that's where I'm most confused

OpenStudy (anonymous):

\[\lim_{x \rightarrow 8}\left( 1\sqrt[3]{x} \right)\left( 2-6x^2+x^3 \right)\] I guess? \lim_{x \rightarrow 8}\left( 1\sqrt[3]{x} \right)\left( 2-6x^2+x^3 \right)

hero (hero):

It's one PLUS the cube root of x

OpenStudy (anonymous):

\[\lim_{x \rightarrow 8}\left( 1+\sqrt[3]{x} \right)\left( 2-6x^2+x^3 \right)\] Weird, very weird...